OpenStudy (anonymous):

Two students are selected at random from a class that consists of 13 males and 7 females. What is the probability that both students are female?

OpenStudy (anonymous):

Alright, listen up, there are 20 students in total, 13 male, 7 female. So if two RANDOM people were chosen, and if female then it would be \[\frac{ 7 }{ 20 }\]Now you take the next student, in this case, one student was chosen (female), and if that happened to be another female you would have \[\frac{ 6 }{ 19 }\]Since that one female left the group, there was then 19 students left for picking and 6 females left. Now take those fractions and multiply them together. \[\frac{ 7 }{ 20 } * \frac{ 6 }{ 19 } = \frac{ 42 }{ 380 }\]

OpenStudy (anonymous):

Simplify your answer down (since math teachers get hard because of that) \[\frac{ 42 }{ 380 } = \frac{ 21 }{ 190 }\]

OpenStudy (anonymous):

Questions?